OpenStudy (anonymous):

Find the perimeter of a square having two consecutive vertices at (7, 2) and (-3, 3). A.4sqrt17 B.4sqrt101 C.20sqrt3 D.4sqrt3

jimthompson5910 (jim_thompson5910):

Find the distance from (7, 2) to (-3, 3)

jimthompson5910 (jim_thompson5910):

Then multiply that distance by 4 to find the perimeter

OpenStudy (anonymous):

i got 1/-10 for distance

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

no, it cfant be right...

jimthompson5910 (jim_thompson5910):

no, the distance has to have a square root in it

OpenStudy (anonymous):

well, how do i find that? i'm sorry if i seem a little slow, but i failed a class last year, and now i have to do it over the computer, so i've been through a whole year in about a week

OpenStudy (anonymous):

but luckily, im already at 90%(: thats why i ask so many questions, this last workseet has 80 problems on it

jimthompson5910 (jim_thompson5910):

I'll show you how to find the distance between these two points and hopefully it will be useful for finding the distance between any two points (so I recommend copying this into your notes) d = sqrt((x2-x1)^2+(y2-y1)^2) d = sqrt((-3-7)^2+( 3- 2)^2) d = sqrt((-10)^2+(1)^2) d = sqrt(100+1) d = sqrt(101) So exact distance between the two points is sqrt(101) units.

OpenStudy (anonymous):

and i dont wanna rush, but i only have half an hour left, and bout 50 questions, so asap please & yeah i've been taking lots of notes..

jimthompson5910 (jim_thompson5910):

so if the distance is sqrt(101), what's the perimeter?

OpenStudy (anonymous):

4sqrt101?

jimthompson5910 (jim_thompson5910):

you got it